# MuLan: Multimodal-LLM Agent for Progressive Multi-Object Diffusion

### v1

@@ 3.4 Conditional Single-Object Diffusion with LLM Planning and Attention Guidance
Figure 3: Single object diffusion with LLM planning and attention guidance for $\texttt{obj}_{n}$ (detailed procedure in Algorithm 1).
@@
1:  Input: Object number $n$, sub-prompt $\texttt{p}_{n}$, LLM planner Planner, precise mask $\bm{\tilde{M}}_{n-1}$ (only for $n>1$), latents $\{\bm{z}_{(n-1),(t-1)}\}_{t=1}^{T}$ (only for $n>1$), attention guidance timestep threshold $T^{\prime}$, combination timestep threshold $T^{*}$ (only for $n>1$), diffusion model $\mathcal{D}$.
@@
At stage-$n$, the diffusion model only focuses on generating $\texttt{obj}_{n}$ according to the sub-prompt $\texttt{p}_{n}$ under attention guidance, which is based on a rough mask (i.e., a bounding box) $\bm{M}_{n}$ for $\texttt{obj}_{n}$. $\bm{M}_{n}$ is derived from $\texttt{obj}_{n}$’s relative position and size planned by the LLM, conditioned on the current space availability in the image. When $n>1$, $\bm{M}_{n}$ also depends on the previous object $\texttt{obj}_{n-1}$ and its precise mask $\tilde{\bm{M}}_{n-1}$ in the image generated in stage-($n$-1). The pipeline is given in Figure 3 with the complete procedure listed in Algorithm 1. We will introduce it step by step in the following.
LLM Planning of a Rough Mask for $\texttt{obj}_{n}$.
[…]
When $n=1$, since there is not any object generated yet, both the position $\texttt{opt}_{1}$ and $\texttt{Num}_{1}$ are unrestricted and the LLM can be prompted to determine $\texttt{opt}_{1}$ and $\texttt{Num}_{1}$ given sub-prompt $\texttt{p}_{1}$. Since the object order starts from left to right and bottom to top, there will be only two position options $\texttt{opt}_{1}\in\{\texttt{left},\texttt{bottom}\}$ for $\texttt{obj}_{1}$. Once $\texttt{opt}_{1}$ determined, as shown in Figure 5, MuLan evenly splits the whole image’s width $W$ ($H$) to $\texttt{Num}_{1}$ parts and assigns the very left (bottom) part to $\texttt{obj}_{1}$, which leads to the following bounding box:
$\bm{M}_{1}=\begin{cases}(0,0,\frac{W}{\texttt{Num}_{1}},H),&\texttt{opt}_{1}=\texttt{left},\\ (\frac{(\texttt{Num}_{1}-1)\cdot H}{\texttt{Num}_{1}},0,W,\frac{H}{\texttt{Num}_{1}}),&\texttt{opt}_{1}=\texttt{bottom}.\end{cases}$ (3)
[…]
$\bm{M}_{n}=\begin{cases}(\tilde{x}_{n-1}+\tilde{w}_{n-1},0,\frac{W-\tilde{x}_{n-1}+\tilde{w}_{n-1}}{\texttt{Num}_{n}},H),\\ \text{if }\texttt{opt}_{n}=\texttt{right},\\ \\ (0,\frac{\tilde{y}_{n-1}\cdot(\texttt{Num}_{n}-1)}{\texttt{Num}_{n}},W,\frac{\tilde{y}_{n-1}}{\texttt{Num}_{n}}),\\ \text{if }\texttt{opt}^{n}=\texttt{top}.\end{cases}$ (4)
[…]
Single-Object Diffusion with Attention Guidance.
Given the rough mask $\bm{M}_{n}$ of $\texttt{obj}_{n}$ in the image, MuLan can precisely generate $\texttt{obj}_{n}$ in the image using the backward guidance method (Chen et al., 2024), which guides the attention masks of the object to be positioned correctly during generation. Specifically, let $B$ be the set of the blocks containing cross-attentions layers in the diffusion model, $\bm{A}^{(j)}$ be the cross-attention map in the $j$-th block, $j\in B$, $\bm{A}^{(j)}_{m,k}$ represents the relevance between the spatial location $m$ and token-$k$ that describes $\texttt{obj}_{n}$ in the prompt, and $|\bm{A}^{(j)}_{(:,k)}|=W\cdot H$ is the cross-attention map of token $k$. Larger values in $\bm{A}^{(j)}_{m,k}$ indicate that $\texttt{obj}_{n}$ is more likely located at the spatial location of $m$. Hence, minimizing the following energy function encourages the cross-attention maps to the $k$-th token ($\texttt{obj}_{n}$) to have a larger value inside the mask:
$\scriptstyle E(\bm{A}^{(j)},\bm{M}_{n},k)=\left(1-\frac{\sum_{m=x_{n}+(y_{n}-1)W}^{x_{n}+w_{n}+(y_{n}+h_{n}-1)W}\bm{A}^{(j)}_{m,k}}{\sum_{m=1}^{W\cdot H}\bm{A}^{(j)}_{m,k}}\right)^{2}.$ (5)
Then, in every step-$t$ of the earlier generation process ($t>T^{\prime}$), MuLan applies gradient descent to minimize the energy and updates the input latent $\bm{z}_{n,t}$ for object $\texttt{obj}_{n}$ by
$\bm{z}_{n,t}=\bm{z}_{n,t}-\eta\cdot\nabla_{\bm{z}_{n,t}}\sum_{j\in B}E(\bm{A}^{(j)},\bm{M}_{n},k),$ (6)
where $\eta$ is the learning rate. In practice, to take the previous objects and their constraints into account when generating $\texttt{obj}_{n}$, during earlier steps of backward diffusion (i.e., $t>T^{*}$), we further combine the latent of $\texttt{obj}_{n}$ and $\texttt{obj}_{n-1}$. Specifically, after step-$t$ of backward diffusion ($t$ varies from $T$ to 00), we update the latent $\bm{z}_{n,(t-1)}$ by
MuLan applies the above single-object diffusion to each object one after another from $\texttt{obj}_{1}$ to $\texttt{obj}_{N}$, as planned by the LLM at the very beginning. The procedure of generating $\texttt{obj}_{n}$ is detailed in Algorithm 1. In each stage-$n$, MuLan applies an LLM planner to determine the location $\texttt{opt}_{n}$ of $\texttt{obj}_{n}$ and object count $\texttt{Num}_{n}$ at $\texttt{opt}_{n}$. MuLan then derives a rough mask $\bm{M}_{n}$ from $\texttt{opt}_{n}$, $\texttt{Num}_{n}$, and previously generated object’s precise mask. Given $\bm{M}_{n}$, MuLan applies a diffusion model to generate $\texttt{obj}_{n}$ in the $\bm{M}_{n}$-defined bounding box via attention guidance.
Overlapping between Objects
Overlapping between objects is a key challenge when generating one object conditioned on the previous one(s). However, it lacks attention in previous methods (Lian et al., 2023; Feng et al., 2023). Instead, we propose an effective strategy that can be merged into the procedure above. Specifically, at the generation of object $\texttt{obj}_{n}$, we prompt the LLM to judge if there is overlapping between $\texttt{obj}_{n}$ and $\texttt{obj}_{n-1}$. If there is no overlapping, the generation is the same as described above. If there is overlapping, we first compute three candidates for the rough mask $\{\bm{M}_{n,i}\}_{i=1}^{3}$, associated with three overlapping ratios $\{r_{i}\}_{i=1}^{3}=\{10\%,30\%,50\%\}$ between $\texttt{obj}_{n-1}$ and $\texttt{obj}^{n}$. An illustration is given in Figure 7 with more details of candidate masks in Appendix C.
Given the three masks $\bm{M}_{n,i}$, we generate three candidate images using Algoirthm 1. Then the CLIP scores (Hessel et al., 2021) between the generated images and the input prompt $\texttt{p}_{n}$ are computed and the image with the maximal CLIP score is selected as the generated image for $\texttt{obj}_{n}$.
3.5 Adaptive Feedback Control by VLM
To correct the possible mistakes made in the sequential generation process, MuLan builds a feedback-loop control by a vision-language model (VLM). After each generation stage, MuLan queries the VLM to inspect the generated object(s)and its consistency with the input prompt. If they do not align well, MuLan will adjust the backward guidance and $T^{*}$ of the current stage to re-generate the object. Such a close-loop control involves LLM, diffusion, and VLM and significantly automates the T2I generation for complicated prompts, leading to a more accurate generation in practice.
4 Experiments
Structure Diffusion (Feng et al., 2022) 88.97%/87.37% 54.62%/62.63% 34.36%/24.24% 64.31%/64.85%
Promptist-SD v1.4 (Hao et al., 2022) 80.36%/70.71% 49.23%/52.02% 24.49%/13.13% 56.73%/51.72%
Promptist-SDXL (Hao et al., 2022) 94.36%/93.94% 70.00%/78.28% 35.89%/33.33% 72.92%/75.56%
SD v1.4 (Rombach et al., 2022) 90.31%/74.49% 57.14%/51.02% 37.24%/32.65% 66.43%/56.73%
SDXL (Podell et al., 2023) 94.64%/78.57% 66.07%/53.06% 41.14%/24.49% 72.34%/57.55%
PixArt-$\alpha$ (Chen et al., 2023) 92.09%/76.53% 66.58%/61.22% 34.69%/32.65% 70.41%/61.63%
Dataset
To evaluate our framework, we construct a prompt dataset from different benchmarks. Specifically, since our focus is to achieve better generation for complex prompts containing multi-objects with both spatial relationships and attribute bindings, we first collect all complex spatial prompts from T2I-CompBench (Huang et al., 2023). To make the experiments more comprehensive, we let ChatGPT generate about 400 prompts with different objects, spatial relationships, and attribute bindings so that the prompt sets consists of about 600 prompts. To further evaluate the capability of our framework on extremely complex and hard prompts, we manually add prompts that SDXL fails to generate, leading to a hard prompt dataset containing 200 prompts. Similar to the complex spatial prompts in T2I-CompBench (Huang et al., 2023), each prompt in our curated dataset typically contains two objects with various spatial relationships, with each object containing attribute bindings randomly selected from {color,shape,texture}.
Models & Baseline
As a training-free framework, MuLan can be incorporated into any existing diffusion models. We evaluate two stable diffusion models with our framework, Stable Diffusion v1.4 (Rombach et al., 2022) and the SOTA Stable Diffusion XL (Podell et al., 2023). To verify the superiority of MuLan, we compare it with previous controllable generation methods and general T2I generation methods. Specifically, we evaluate Structure Diffusion (Feng et al., 2022), Promptist (Hao et al., 2022), the original Stable Diffusion v1.4, the original SDXL, and the recent SOTA diffusion model PixArt-$\alpha$ (Chen et al., 2023).
Implementation Details
MuLan use GPT-4 (Achiam et al., 2023) as the LLM planner, and LLaVA-1.5 (Liu et al., 2023) as the VLM checker to provide the feedback. We also conducted an ablation study to show the importance of the feedback control provided by the VLM and the effect of different VLMs. Moreover, we found the attention blocks utilized during the attention guidance are vital, which can be classified as near-input blocks, near-middle blocks, and near-output blocks. We utilize the near-middle blocks in our main experiments and also show the ablation results of different blocks.
Evaluation
Since the prompt dataset contains texts with complex compositions, we design a questionnaire to comprehensively investigate the alignment between the generated image and the corresponding input text. The questionnaire is composed of three aspects - object completeness, correctness of attribute bindings, and correctness of spatial relationships. We only set two options for each question (Yes or No), without any ambiguity. For detailed questions and examples of the evaluation, please refer to Appendix D. For each aspect of the evaluation, we compute the percentage of answers with “Yes”. Given the generated image, we assess the image’s quality using a questionnaire asking both the state-of-the-art multi-modal large language model (GPT-4V (OpenAI, 2023)) and the human evaluator.
4.1 Main Results and Analysis
Results with GPT Evaluation
Given the generated image, we prompt GPT-4V to answer the questions about the image in the questionnaire, where each only focuses on one of the three aspects. The results for different methods and different base models are shown in Table 1. The results show that our framework can achieve the best performance compared to different controllable generation methods and T2I generation methods. In particular, in the two ‘harder’ aspects - attribute bindings and spatial relationships, MuLan can surpass other methods by a large margin. More qualitative results can be found in Figure 6 and Appendix E.
Results with Human evaluation
4.2 Ablation study
Ablation on the attention blocks
Table 2: Ablation study on attention blocks with SD-v1.4 as the base model. “Objects”, “Attributes”, and “Spatial” denote Object completeness, Attribute bindings, and Spatial relationships. The results (evaluated by GPT-4V (OpenAI, 2023)) show that near-middle attention blocks perform the best for attention guidance.
Guidance Objects Attributes Spatial Overall
near-input 83.67% 55.10% 14.29% 58.37%
near-middle 97.96% 80.61% 30.61% 77.55%
near-output 72.45% 45.92% 22.45% 51.84%
Ablation on the VLM feedback control
The VLM feedback control is a key componenet in MuLan to provide feedback and adjust the generation process to ensure the every stage’s correct generation. Here, we show the importance of the feedback by removing feedback control from the whole framework. As shown in Table 3, after removing the VLM, the results would be much worse. It is because there is no guarantee or adaptive adjustment for each generation stage, which verifies that the feedback control provided by the VLM is essential to handle complex prompts. Moreover, we also test MuLan’s compatibility with different VLMs. As shown in Table 4, we compare the Mulan’s performance using different VLMs including LLaVA-1.5 (Liu et al., 2023), GPT-4V (OpenAI, 2023), and Gemini-Pro (Team et al., 2023). The results show that MuLan could still maintain a good performance with different choices of the VLM and achieve good compatibility.
MuLan Objects Attributes Spatial Overall
w/ Feedback 97.96% 80.61% 30.61% 77.55%
w/o Feedback 81.63% 59.18% 18.37% 60.00%
VLM in MuLan Objects Attributes Spatial Overall
LLaVA-1.5 (Liu et al., 2023) 97.96% 80.61% 30.61% 77.55%
GPT-4V (OpenAI, 2023) 95.92% 80.61% 28.57% 76.33%
Gemini-Pro (Team et al., 2023) 95.92% 83.67% 38.78% 79.59%
5 Conclusions and Limitations
In this paper, we propose a training-free multimodal-LLM agent (MuLan) to progressively generate objects contained in the complicated input prompt with closed-loop feedback control, achieving better and more precise control on the whole generation process. By first decomposing the complicated prompt into easier sub-tasks, our method takes turns to deal with each object, conditioned on the previous one. The VLM checker further provides a guarantee with feedback control and adaptive adjustment for correct generation at each stage. Extensive experiments demonstrate the superiority of MuLan over previous methods, showing the potential of MuLan as a new paradigm of controllable diffusion generation. However, there are still limitations to be further addressed in the future work. Since the whole generation contains multiple stages, depending on the number of objects, it will take a longer time than a one-stage generation approach. On the other hand, the LLM planner may mistakenly parse the input prompt which results in incorrect decomposition. This could be addressed by first re-writing the input prompt by the LLM to facilitate later processing.
Broader Impact
Appendix A Detailed prompt template of the global planning by the LLM

### v2

@@ 3.3 Conditional Single-Object Diffusion with LLM Planning and Attention Guidance
Figure 3: Single object diffusion with LLM planning and attention guidance for $\texttt{obj}_{n}$ (detailed procedure in Algorithm 1 in Appendix D).
LLM Planning of a Rough Mask for $\texttt{obj}_{n}$.
[…]
Once $\bm{M}_{n}$ is determined, the cross-attention guidance is utilized during generation of $\texttt{obj}_{n}$ to ensure $\texttt{obj}_{n}$ is correctly generated within $\bm{M}_{n}$, as elaborated in the following.
Single-Object Generation with Attention Guidance.
Given the rough mask $\bm{M}_{n}$ of $\texttt{obj}_{n}$, the next is to ensure the generated $\texttt{obj}_{n}$ will be correctly located within $\bm{M}_{n}$. A natural and intuitive way to achieve this in diffusion models is to guide the generation of the cross-attention map of $\texttt{obj}_{n}$, which builds the relevance between the text prompt and the location of generated object.
To this end, MuLan manipulates the cross-attention map of $\texttt{obj}_{n}$ under the guidance of $\bm{M}_{n}$, using the backward guidance method [4], to maximize the relevance inside $\bm{M}_{n}$. Specifically, let $\bm{A}$ be the cross-attention map, $\bm{A}_{m,k}$ represents the relevance between the spatial location $m$ and token-$k$ that describes $\texttt{obj}_{n}$ in the prompt. Larger value in $\bm{A}_{m,k}$ indicates that $\texttt{obj}_{n}$ is more likely located at the spatial location of $m$. The goal is to maximize the relevance $\bm{A}_{m,k}$ inside the mask $\bm{M}_{n}$ while minimizing the relevance outside the mask $\bm{M}_{n}$. Hence the following energy function is utilized:
$\scriptstyle E(\bm{A},\bm{M}_{n},k)=\left(1-\frac{\sum_{m\in\bm{M}_{n}}\bm{A}_{m,k}}{\sum_{m}\bm{A}_{m,k}}\right)^{2},$ (1)
where $\sum_{m\in\bm{M}_{n}}$ denotes the summation over the spatial locations included in $\bm{M}_{n}$, and $\sum_{m}$ denotes the summation over all the spatial locations in the attention map. In every step-$t$ of the earlier generation process, MuLan applies gradient descent to minimize the energy by updating the input latent $\bm{z}_{n,t}$ for object $\texttt{obj}_{n}$. In this way, the cross-attention map corresponding to $\texttt{obj}_{n}$ will achieve the largest relevance inside $\bm{M}_{n}$, meaning $\texttt{obj}_{n}$ can be correctly positioned inside the rough mask.
On the other hand, to take the previous objects and their constraints into account when generating $\texttt{obj}_{n}$, we further combine the latent of $\texttt{obj}_{n}$ and $\texttt{obj}_{n-1}$. Specifically, after step-$t$ of reverse process ($t$ varies from $T$ to 00), we update the latent $\bm{z}_{n,(t-1)}$ by
MuLan applies the above single-object diffusion to each object one after another from $\texttt{obj}_{1}$ to $\texttt{obj}_{N}$, as planned by the LLM at the very beginning. The procedure of generating $\texttt{obj}_{n}$ is detailed in Algorithm 1.
Objects Overlapping.
Overlapping between objects is a key challenge in text-to-image diffusion models. However, it lacks attention in previous methods [12, 6]. Instead, we propose an effective strategy that can be merged into the procedure above. Specifically, at the generation of object $\texttt{obj}_{n}$, we prompt the LLM to judge if there is overlapping between $\texttt{obj}_{n}$ and $\texttt{obj}_{n-1}$. If there is overlapping, we first compute three candidates for the rough mask $\{\bm{M}_{n,i}\}_{i=1}^{3}$, associated with three overlapping ratios $\{r_{i}\}_{i=1}^{3}=\{10\%,30\%,50\%\}$ between $\texttt{obj}_{n-1}$ and $\texttt{obj}_{n}$.
Given the three masks $\bm{M}_{n,i}$, MuLan generates three candidate images using Algorithm 1. Then the CLIP scores [8] between the generated images and the input prompt $\texttt{p}_{n}$ are computed and the image with the maximal CLIP score is selected as the generated image for $\texttt{obj}_{n}$. An illustration is given in Figure 8 with more details of candidate masks in Appendix H.
3.4 Interaction with VLM and Human Users during Generation
To correct the possible mistakes made in the sequential generation process, MuLan builds an adaptive feedback-loop control by interacting with a vision-language model (VLM). After each generation stage, MuLan queries the VLM to inspect the generated object(s) and its consistency with the input prompt. If they do not align well, MuLan will adjust the backward guidance of the current stage to re-generate the object. Such a close-loop control involves LLM, diffusion, and VLM and significantly automates the T2I generation for complicated prompts, leading to a more accurate generation in practice.
Figure 4: An illustration tree for difference cases of human-agent interaction during generation. The middle branch (connected by blue arrows) shows the original generation process without human-agent interaction. The top and bottom branches show different complex composed human-agent interaction during generation for various adjustments, involving object adjustments, attribute adjustments, and spatial relationship adjustments, which demonstrate the flexibility and effectiveness of MuLan for human-agent interaction during generation.
In addition, the multi-step process naturally allows human-agent interaction/collaboration during generation in practice. Users can timely monitor the generation process. In this way, the interaction enables users to make preferred changes and adjustments to the generated images easily and effectively by providing adjusting prompts to MuLan at any intermediate step, such as attribute adjustment, object adjustment, and spatial relationship adjustment. With the adjusting prompts, MuLan will utilize the LLM to modify the original prompt accordingly and change the generation process to the preferred one. An illustration for different changes or adjustments during generation is shown in Figure 4, which indicates MuLan can achieve both simple and composed complex adjustments with interaction. In contrast, for other existing generation and editing methods, users have to wait until the whole generation process is finished. Therefore, the proposed framework is more user-friendly and flexible in terms of human-agent interaction and collaboration.
4 Experiments
Dataset
To evaluate our framework, we construct a prompt dataset from different benchmarks. Specifically, since our focus is to achieve better generation for complex prompts containing multi-objects with both spatial relationships and attribute bindings, we first collect all complex spatial prompts from T2I-CompBench [10]. To make the experiments more comprehensive, we let ChatGPT generate about 400 prompts with different objects, spatial relationships, and attribute bindings so that the prompt sets consists of about 600 prompts. To further evaluate the capability of our framework on extremely complex and hard prompts, we manually add prompts that SDXL fails to generate, leading to a hard prompt dataset containing 200 prompts. Similar to the complex spatial prompts in T2I-CompBench [10], each prompt in our curated dataset typically contains two objects with various spatial relationships, with each object containing attribute bindings randomly selected from {color,shape,texture}.
Models & Baseline
As a training-free framework, MuLan can be incorporated into any existing diffusion models. We evaluate two stable diffusion models with our framework, Stable Diffusion v1.4 [17] and the SOTA Stable Diffusion XL [16]. To verify the superiority of MuLan, we compare it with previous controllable generation methods and general T2I generation methods. Specifically, we evaluate Structure Diffusion [5], Promptist [7], the original Stable Diffusion v1.4, the original SDXL, and the recent SOTA diffusion model PixArt-$\alpha$ [3].
Implementation Details
MuLan use GPT-4 [1] as the LLM planner, and LLaVA-1.5 [13] as the VLM checker to provide the feedback. We also conducted an ablation study to show the importance of the feedback control provided by the VLM and the effect of different VLMs. Moreover, we found the attention blocks utilized during the attention guidance are vital, which can be classified as near-input blocks, near-middle blocks, and near-output blocks. We utilize the near-middle blocks in our main experiments and also show the ablation results of different block. All the experiments are conducted on a single NVIDIA RTX A6000 GPU.
Evaluation
Since the prompt dataset contains texts with complex compositions, we design a questionnaire to comprehensively investigate the alignment between the generated image and the corresponding input text. The questionnaire is composed of three aspects - object completeness, correctness of attribute bindings, and correctness of spatial relationships. We only set two options for each question (Yes or No), without any ambiguity. For detailed questions and examples of the evaluation, please refer to Appendix I. For each aspect of the evaluation, we compute the percentage of answers with “Yes”. Given the generated image, we assess the image’s quality using a questionnaire asking both the state-of-the-art multi-modal large language model (GPT-4V [15]) and the human evaluator.
4.1 Main Results and Analysis
Results on GPT Evaluation
Given the generated image, we prompt GPT-4V to answer the questions about the image in the questionnaire, where each only focuses on one of the three aspects. The results for different methods and different base models are shown in Table 1. The results show that our framework can achieve the best performance compared to different controllable generation methods and T2I generation methods. In particular, in the two ‘harder’ aspects - attribute bindings and spatial relationships, MuLan can surpass other methods by a large margin. More qualitative results can be found in Figure 5 and Appendix J.
Structure Diffusion [5] 88.97%/87.37% 54.62%/62.63% 34.36%/24.24% 64.31%/64.85%
Promptist-SD v1.4 [7] 80.36%/70.71% 49.23%/52.02% 24.49%/13.13% 56.73%/51.72%
Promptist-SDXL [7] 94.36%/93.94% 70.00%/78.28% 35.89%/33.33% 72.92%/75.56%
SD v1.4 [17] 90.31%/74.49% 57.14%/51.02% 37.24%/32.65% 66.43%/56.73%
SDXL [16] 94.64%/78.57% 66.07%/53.06% 41.14%/24.49% 72.34%/57.55%
PixArt-$\alpha$ [3] 92.09%/76.53% 66.58%/61.22% 34.69%/32.65% 70.41%/61.63%
Results on Human Evaluation
Results on Human-Agent Interaction
To show MuLan is still very effective if users want to modify the input prompt or edit the generated images during the generation, i.e., the human-agent interaction, we use ChatGPT to mimic the user to generate various adjusting prompts for the interaction with MuLan on randomly sampled 50 prompts. SD v1.4 [17] is utilized as the base model. The generated adjusting prompts focus on several aspect, i.e., attribute adjustment, object adjustment, and spatial relationship adjustment. We use GPT-4V [15] to quantitatively evaluate the performance of MuLan given the final generated images and final text prompts, as shown in Table 2. The results indicate that MuLan can still achieve high accuracy even with various adjustments/changes during generation.
Table 2: GPT-4V evaluation of final generated images and final prompts after adjustments/changes. The results show that MuLan is still very effective with various adjustment of prompts during generation.
Objects Attributes Spatial Overall
MuLan-SD v1.4 95.92% 72.45% 28.57% 73.06%
4.2 Ablation study
Ablation on the attention blocks
Table 3: Ablation study on attention blocks with SD-v1.4 as the base model. “Objects”, “Attributes”, and “Spatial” denote Object completeness, Attribute bindings, and Spatial relationships. The results (evaluated by GPT-4V [15]) show that near-middle attention blocks perform the best for attention guidance.
Guidance Objects Attributes Spatial Overall
near-input 83.67% 55.10% 14.29% 58.37%
near-middle 97.96% 80.61% 30.61% 77.55%
near-output 72.45% 45.92% 22.45% 51.84%
Ablation on the VLM feedback control
The VLM feedback control is a key componenet in MuLan to provide feedback and adjust the generation process to ensure the every stage’s correct generation. Here, we show the importance of the feedback by removing feedback control from the whole framework. As shown in Table 4, after removing the VLM, the results would be much worse. It is because there is no guarantee or adaptive adjustment for each generation stage, which verifies that the feedback control provided by the VLM is essential to handle complex prompts. Moreover, we also test MuLan’s compatibility with different VLMs. As shown in Table 5, we compare the Mulan’s performance using different VLMs including LLaVA-1.5 [13], GPT-4V [15], and Gemini-Pro [21]. The results show that MuLan could still maintain a good performance with different choices of the VLM and achieve good compatibility.
MuLan Objects Attributes Spatial Overall
w/ Feedback 97.96% 80.61% 30.61% 77.55%
w/o Feedback 81.63% 59.18% 18.37% 60.00%
VLM in MuLan Objects Attributes Spatial Overall
LLaVA-1.5 [13] 97.96% 80.61% 30.61% 77.55%
GPT-4V [15] 95.92% 80.61% 28.57% 76.33%
Gemini-Pro [21] 95.92% 83.67% 38.78% 79.59%
5 Conclusions and Limitations
In this paper, we propose a training-free multimodal-LLM agent (MuLan) to progressively generate objects contained in the complicated input prompt with closed-loop feedback control, achieving better and more precise control on the whole generation process. By first decomposing the complicated prompt into easier sub-tasks, our method takes turns to deal with each object, conditioned on the previous one. The VLM checker further provides a guarantee with feedback control and adaptive adjustment for correct generation at each stage. Moreover, the application to the human-agent interaction further enhances the significance of MuLan, making the generation more flexible and effective to align with the preferences of users. Extensive experiments demonstrate the superiority of MuLan over previous methods, showing the potential of MuLan as a new paradigm of controllable diffusion generation. However, there are still limitations to be further addressed in the future work. Since the whole generation contains multiple stages, depending on the number of objects, it will take a longer time than a one-stage generation approach. On the other hand, the LLM planner may mistakenly parse the input prompt which results in incorrect decomposition. This could be addressed by first re-writing the input prompt by the LLM to facilitate later processing.
Appendix A Broader Impact
Appendix B Differences between MuLan and the concurrent work RPG
As stated in Introduction and Related work, although we acknowledge that our proposed framework shares a similar high-level idea with RPG, we would like to emphasize that there are still substantial differences between ours and RPG.
Firstly, our proposed MuLan aims to progressively generate each object given each subprompt. At the same time, the objects are generated conditioned on previously generated objects. In RPG, on the other hand, all objects are generated independently. In addition, different from RPG which requires manually designed in-context examples for the CoT reasoning, ours does not have such requirement. We directly utilize LLMs for the planning during generation, which is an easier task and can be done by LLMs without in-context learning. What’s more, MuLan can adaptively control and correct the generation results using feedback by the VLMs while RPG does not have the feedback for the generation. Also, for the common overlapping problem between objects, we propose a strategy to generate several candidates to deal with it. In contrast, in RPG, the overlapping parts are treated as a whole for generation.
More importantly, as we show in Section 4.1, our proposed framework can be directly applied to human-agent interaction during generation to facilitate flexible and effective changes/adjustments of the process while RPG cannot achieve the interaction. To summarize, the main differences between MuLan and RPG are as follows:
  • Our proposed MuLan generates each object conditioned on previously generated objects while RPG generates all objects in parallel independently.
  • MuLan does not require any in-context learning during the whole generation; in RPG, specifically designed in-context examples are needed for Chain-of-Thought reasoning.
  • MuLan utilizes the VLM-based feedback control to ensure each object can be generated correctly while RPG does NOT have such a feedback mechanism.
  • We propose a strategy to deal with overlapping/interaction between objects whereas RPG directly treats overlapping objects as a whole part to generate.
  • MuLan can be directly applied to human-agent interaction during generation for flexible and various adjustments of the generation process while RPG cannot achieve it.
Appendix C Background on (Latent) Diffusion Models
Consisting of the diffusion process and the reverse process, diffusion models have shown impressive capability for high-quality image generation by iteratively adding noise and denoising [9]. Let $\bm{x}_{0}\sim q(\bm{x}_{0})$ be the true data distribution. Starting from $\bm{x}_{0}$, the diffusion process adds different levels of noise pre-defined by the schedule $\{\beta_{t}\}_{1}^{T}$, producing $\bm{x}_{1},\cdots,\bm{x}_{T}$. As $T\to\infty$, $\bm{x}_{T}$ will become the standard Gaussian distribution $\mathcal{N}(\bm{0},\bm{I})$. Accordingly, the reverse process aims to reverse the above process and reconstruct the true data distribution from $p(\bm{x}_{T})=\mathcal{N}(\bm{0},\bm{I})$ by a parameterized noise model $\epsilon_{\theta}(\cdot)$. With $\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I})$, the training loss of the model can be simplified as
Latent diffusion models [17] have recently attracted growing attention due to their efficiency and superior performance. Instead of performing diffusion and its reverse process in the pixel space, they add noise and denoise in a latent space of $\bm{z}$ encoded by a pre-trained encoder $\mathcal{E}$. Thereby, the diffusion process starts from $\bm{z}_{0}=\mathcal{E}(\bm{x}_{0})$ and subsequently produces latent states $\bm{z}_{1},\cdots,\bm{z}_{t},\cdots,\bm{z}_{T}$. Accordingly, the training loss becomes
Appendix D Algorithm procedure of single-object diffusion in MuLan
The complete and detailed procedure of single object diffusion described in Section 3.3 is shown in Algorithm 1.
@@ Appendix D Algorithm procedure of single-object diffusion in MuLan
1:  Input: Object number $n$, sub-prompt $\texttt{p}_{n}$, LLM planner Planner, precise mask $\bm{\tilde{M}}_{n-1}$ (only for $n>1$), latents $\{\bm{z}_{(n-1),(t-1)}\}_{t=1}^{T}$ (only for $n>1$), attention guidance timestep threshold $T^{\prime}$, combination timestep threshold $T^{*}$ (only for $n>1$), learning rate $\eta$, diffusion model $\mathcal{D}$.
@@
Appendix E Detailed prompt template of the global planning by the LLM